OpenStudy (anonymous):

what is 5log5^3 = ??

OpenStudy (anonymous):

use the following property og log \[\Large \log b^n=nlog (b)\] can you do this now?

OpenStudy (anonymous):

@king this property say the exponent gets down and gets multiplied with log so 3 should come down and multiplied with 5 can you do this ?

OpenStudy (theviper):

\[\Large{5\log_53=3}\]because\[\Huge{alog_ab=b}\]

OpenStudy (anonymous):

@TheViper do not you think question is?? \[\Huge 5\log5^3\] where log is base 10.

OpenStudy (anonymous):

10^3

OpenStudy (theviper):

Can u rewrite the full question @sami-21 ??

OpenStudy (anonymous):

what is 5log5^3 =

OpenStudy (theviper):

\[\Large{5\log_53\space ??}\]

OpenStudy (theviper):

If the question then I m right :P

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

it could also be 5^log^3=